1. Assume that in 2015, the following prevails in the republic of Nurd

Y= $200

C= $160

S= $40

I (Planned)= $30

G= $0

T= $0


Assume that households consume 80 percent of their income, they save 20 percent of their income, MPC = 0.8, and MPS = 0.2. That is, C = 0.8Yd and S = 0.2Yd.


a. Is the economy of Nurd in equilibrium? What is Nurd’s equilibrium level of income? What is likely to happen in the coming months if the government takes no action?


b. If $200 is the “full-employment” level of Y, what fis-cal policy might the government follow if its goal is full employment? c. If the full-employment level of Y is $250, what fiscal policy might the government follow?


c. If the full-employment level of Y is $250, what fiscal policy might the government follow?




1
Expert's answer
2021-12-13T16:52:44-0500

a)

C + I + G + Xn = C + S+ T.


160+30=160+40

190=200


the economy is out of balance equilibrium

 equilibrium 200

When the level of actual GDP is below the potential level, aggregate demand is less than aggregate potential supply. There will be unemployment and there will be a downward trend in the price level in the economy in a deflationary environment

b)Stimulating fiscal policy is applied

c)Stimulating fiscal policy is applied